# Joint Pruning and Channel-wise Mixed-Precision Quantization for Efficient Deep Neural Networks

### v1

@@ Abstract
Deep Learning, Edge Computing, Quantization, Pruning, NAS
@@ 1 Introduction
Deep neural networks (DNNs) have showcased impressive performance in a wide range of domains; nevertheless, their computational complexity often clashes with the resource limitations of hardware (HW) devices, especially at the edge [1].
Consequently, several studies have explored techniques aimed at reducing the computational complexity and memory requirements of DNNs, while preserving their task performance. Pruning[2] and quantization[3] are two of the most popular ones. The former eliminates unimportant computations from a network, reducing the number of parameters and operations, while quantization, in particular with Mixed-Precision Search (MPS), optimizes the data representation of the model’s parameters and activations.
@@
Our code is open-source at: https://github.com/eml-eda/mixprec-pruning The paper is structured as follows. Section 2 covers the required background concepts, and Section 3 summarizes the most relevant works on mixed-precision quantization and pruning. In Section 4 we illustrate our proposed approach and in Section 5 the experimental results. Lastly, Section 6 concludes the paper.
@@ 2.1 Quantization and Mixed-Precision Search
Quantization is one of the key optimizations to reduce the complexity of DNNs. In particular, integer quantization is extensively used to replace the floating-point representation of both weights and activations with low bit-width integers. On the one hand, low bit-width integers consume less memory compared to floating-point, resulting in model compression. On the other, integer arithmetic operators are faster and more energy-efficient [3]. Moreover, integer quantization also enables the execution of DNNs even on edge devices without a Floating Point Unit (FPU). Although quantization can be applied on networks already trained in float (i.e., post-training quantization), simulating its effect during training (i.e., quantization-aware training, or QAT) is often beneficial for accuracy [3].
@@
Conventional quantization uses fixed-precision, wherein although different parts of a DNN can use different quantization parameters, a uniform bit-width, typically 8 bits, is applied throughout the entire model. Recently, however, mixed-precision strategies have gained attention [7]. They involve using different bit-widths for different sections of a DNN, resulting in additional optimization opportunities in terms of time, memory, and energy consumption, especially when the underlying hardware natively supports operations at sub-byte precision [9, 10]. Nonetheless, determining the optimal allocation of bit-widths to different segments of the network presents a substantial challenge, as it entails exploring a large solution space that grows exponentially with the DNN’s depth. Recently, approaches inspired by Neural Architecture Search (NAS) [16, 17] have been applied to solve this so-called Mixed-Precision Search (MPS) problem. Such approaches can be classified into two families, namely iterative and one-shot.
@@ 2.2 Pruning
The granularity identifies the two main pruning approaches, i.e., unstructured and structured. Unstructured pruning considers the removal of single weights. This approach leads to the largest theoretical reduction of parameters and operations without accuracy loss, reaching values as high as 90% [22]. However, it leads to sparsely connected networks that cannot be easily accelerated in hardware, especially on general-purpose platforms, without adding a significant overhead [23, 21]. In fact, sparse computations make it complex to achieve high utilization on parallel hardware and damage the regularity and locality of memory access patterns, resulting in worse cache behaviour [23].
@@ 3.1.1 Non-differentiable techniques
Several recent works have been devoted to the exploration of automatic methods to optimally assign different precisions to different parts of DNNs. The first attempts trying to solve the MPS problem considered only task performance as optimization goal. HAWQ-V2 [34] uses a sensitivity-based heuristic approach that considers second-order Hessian information. Lin et al. [35] show an analytical method that optimizes signal-to-quantization-noise-ratio (SQNR) to find the optimal bit-width allocations across the layers of the network. HAQ [18] and ReLeQ [19] propose to use an iterative scheme, where an RL agent drives the bit-width assignment, at the granularity of individual layers, using both task performance and latency or energy measurements from the target hardware as rewards. Such approaches, while effective and flexible, suffer from the problem of long convergence times due to the iterative search space exploration, as discussed in Sec. 2.1.
@@ 3.1.2 Differentiable techniques
One-shot algorithms emerged as a more lightweight solution with respect to the non-differentiable black-box optimization techniques. The earliest methods were based on a supernet scheme, inspired by the Differentiable NAS literature [17]. A multi-path network (the supernet) is constructed, which, in the case of MPS, contains all possible bit-width assignments as alternative paths. The optimization goal then reduces to selecting a single path from this network, and is solved through continuous relaxation using gradient-descent. Wu et al. [27] propose one of the first approaches of this kind, where the optimization is performed considering a cost penalty term associated to the memory along with the standard task-specific loss. Gong et al. [30] propose a similar approach where precisions are explored on MobileNet-V2 architectures using as cost the energy consumption on the BitFusion accelerator.
One of the key limitations of supernet-based MPS is the linear scaling of memory and computation complexity with the size of the search space. That is, adding a new path to the supernet with an operator quantized at a different precision requires duplicating the weights and computations associated with such an operator. For this reason, the search can be performed only on simpler and smaller proxy tasks, severely limiting the applicability of the method. EdMIPS [7] solves this problem by substituting the expensive parallel convolutions of supernet-like approaches with an efficient composite convolution operation. In this approach, a single copy of the float tensors is fake-quantized on-the-fly at the different precisions explored (e.g., 2-, 4- and 8-bit) in each iteration of the gradient-based search.
@@ 3.2.2 Differentiable techniques
DJPQ [32] proposes a joint pruning and quantization scheme where quantization is non-linear and pruning is based on Gaussian gates that control the weight distribution of each layer’s channel. They learn a variational posterior that tries to minimize the mutual information between successive layers’ outputs. Bayesian-bits [31] presents a holistic framework for structured channel pruning and mixed-precision quantization, where pruning is considered as “0-bit” precision. The optimization uses learnable stochastic gates, each associated with a specific precision, and is performed to balance the task performance and network complexity, measured as bitops. One limitation of this approach is that it is structurally able to explore only precisions which are power-of-two. This can be a drawback for HW platforms such as NE16, which support weights quantized to all possible bit-widths from 2 to 8. Note that this limitation is the mirror image of the one highlighted for FracBits [28]. Moreover, both DJPQ [32] and Bayesian Bits [31] consider a per-layer mixed-precision scheme and rely on cost metrics that poorly correlate with real HW quantities, such as bitops [37].
Chitty-Venkata et al. [33] propose a joint MPS and pruning approach based on gradient-descent. However, quantization is only performed layer-wise and pruning uses a 2:4 block-balanced approach with a fixed sparsity of 50%. This limits the benefits of this approach to hardware platforms that support block-sparse operations. Chitty-Venkata et al. [39] similarly apply layer-wise MPS with a differentiable algorithm, proposing a method to eliminate sub-optimal configurations from the search space.
@@ 4.2 Optimization Method
where $W$ is the set of network’s weights, and $\theta:=\{\delta,\gamma\}$ is the set of bit-width selection parameters. The task loss term $\mathcal{L}_{\text{task}}(W,\theta)$ depends on both $W$ and $\theta$. On the other hand, the regularization loss term $\mathcal{R}(\theta)$ is only affected by the bit-width selection parameters, and must be designed to model the desired cost metric, such as size or latency, in a differentiable manner. The relative balance between the two terms is controlled by the scalar strength hyper-parameter $\lambda$. A higher value of $\lambda$ biases the search process towards finding more light-weight networks.
In the search phase, for each weight channel or activation tensor of supported layers (convolutional and linear), an effective tensor is computed, as represented in Figure 2. More in detail, considering an activation tensor $X^{(n)}$ and its associated bit-width selection parameters $\delta^{(n)}$, the first step consists in obtaining a discrete probability distribution from the latter. Thus, we obtain a new vector $\hat{\delta}^{(n)}$ whose elements are constrained in the $[0,1]$ range and sum up to 1. This can be achieved by applying one among different sampling methods. In this paper we consider three sampling methods, namely a softmax operation (SM), an argmax (AM) operation and a hard Gumbel-Softmax operation (HGSM). We can thus define the sampling operation for the bit-width selection parameters as $h(\delta)$, such that for $i=0,...,|P_{X}|$:
@@ 4.3 Complexity Regularizers
The methodology detailed in Sections 4.1 and 4.2 is orthogonal to the cost metric selected as proxy for the DNN’s complexity ($\mathcal{R}$ in Eq. 2). Thus, any measure of DNN cost can be applied to guide the training process, as long as it is expressed by a differentiable function, given that it must be included in a gradient-based optimization loop. Commonly used cost functions are hardware-agnostic proxies for general properties of a DNN, like its size or its number of Multiply-and-Accumulate (MAC) operations [40]. While model size is a good proxy for memory occupation, the number of MACs often correlates poorly with the measured performance (e.g., inference latency or energy consumption) on the real hardware [41]. Therefore, more accurate regularizers can be introduced, which take into account a given platform’s support for each combination of weights and activation precision, and the corresponding efficiency.
@@ 4.3.1 Size Regularizer
where $K_{x}^{(n)}$ and $K_{y}^{(n)}$ are the horizontal and the vertical kernel size. The term $C_{\text{in,eff}}^{(n)}$ represents the effective number of input channels, i.e., those which have not been pruned away. Specifically, for a simple case of sequentially-connected layers, $C_{\text{in,eff}}^{(n)}$ is computed as $C_{\text{out}}^{(n-1)}-\sum_{i=1}^{C_{\text{out}}^{(n-1)}}\hat{\gamma}_{i,p_{0}}$. Using $C_{\text{in,eff}}$ instead of $C_{\text{in}}$ in the cost function models the fact that pruning away an output feature map also provides benefits to the subsequent layers because they have to process a smaller input.
@@ 4.3.2 Mixed Precision Inference Core Regularizer
To quantify the latency cost of executing a DNN on MPIC, we use the Look-Up Table (LUT) from [9], which comprises latency measurements collected on fixed layer topologies. In particular, the LUT stores the number of MAC operations per cycle for all combinations of activations and weights bit-widths. For a single convolutional layer $n$, the MPIC regularizer $\mathcal{R}^{(n)}(\hat{\delta}^{(n-1)},\hat{\gamma}^{(n)})$ is defined as:
@@ 4.4 Training Procedure
The procedure to apply the proposed joint mixed-precision and pruning optimization comprises three distinct phases: warmup, search and fine-tuning. The warmup phase consists in the training of the DNN’s weights only, without the additional bit-width selection parameters. Weights are trained in floating point, and optimized only with respect to the task loss, whereas the regularization loss is not applied. The DNN generated by this preliminary training serves as starting point for the subsequent optimization phases.
@@ 5 Experimental Results
Figure 5: Comparison of the results obtained by our proposed method with other state-of-the art approaches. The architecture optimized by PIT that is used as input for MixPrec is denoted as PIT Seed.
@@ 5.1 Experimental Setup
We run experiments on three different benchmarks, namely CIFAR-10  [^3] http://www.cs.toronto.edu/~kriz/cifar.html, Google Speech Commands (GSC) v2 [11] and Tiny ImageNet  [^4] Le, Y., & Yang, X.S. (2015). Tiny ImageNet Visual Recognition Challenge.. For CIFAR-10 we divide the images into training, validation and test sets in the proportions of 66%, 17% and 17%, respectively. The reference architecture is a custom ResNet with 9 convolutional layers, as in [44]. For GSC, as in [11], we consider the standard classification setting comprising only 12 target labels, i.e., 10 keywords from the original dataset and two additional ”Silence” and ”Unknown word” classes. The training, validation and test sets include, respectively, 85%, 10% and 5% of the data samples. The reference architecture is the Depthwise Separable Convolutional Neural Network (DS-CNN) from [44]. For Tiny ImageNet, we divide the original training set into a training and a validation partition with a 90:10 split. Then, we used the official validation partition as test set. The baseline DNN is a ResNet-18. All accuracy results refer to the respective test sets.
@@ 5.2 Sampling Methods Comparison
Fig. 4 reports the results obtained by applying our method on the three considered benchmarks, using the size regularizer as cost metric. Each point in the curves represents a model trained with a different value of the regularization strength ($\lambda$). Only the DNNs belonging to the Pareto front according to the validation metrics are reported. Each plot shows as baselines the reference architecture trained in floating point (FP) and three fixed-precision versions with the weights quantized at 2, 4, and 8 bits. We keep the activations quantized at 8 bits since they do not have an impact on the model size. Each color of the curve is associated to a different sampling method.
@@
Based on these experiments, we conclude that the Softmax sampling strategy is the most stable one, since it yields the best results on CIFAR-10 and GSC, and has a comparable performance with respect to the other sampling methods on Tiny ImageNet. For this reason, we employ Softmax sampling in all following Sections.
@@ 5.3 State-of-the-art Comparison
We present the results on GSC in the central plot of Fig. 5. EdMIPS and MixPrec yield comparable results with each other, with very similar Pareto fronts. Our method achieves an accuracy of 91.60% with a model size of 2.98 kB. At iso-accuracy, we reduce size by 56.17% with respect to the 6.79 kB model found by MixPrec. Our method also leads to comparable results with respect to the concatenation of PIT and MixPrec. This is due to the fact that the reference FP model is highly over-sized, thus the pruning performed by PIT does not remove fundamental channels, which cannot be recovered afterwards. Furthermore, quantization at low bit-widths does not wreck accuracy, allowing MixPrec to reduce the precision with negligible performance impact.
However, as shown in Table II our method significantly reduces the total search time with respect to the concatenation of PIT and MixPrec. Individually, a single PIT epoch requires 1.8×1.8$\times$1.8 × more time than a standard training epoch of the reference DNN, while for both MixPrec and our proposed approach, the overhead is 4.3×4.3$\times$4.3 ×. On the other hand, obtaining a seed for MixPrec when the two methods are applied sequentially requires to first derive the full Pareto curves of architectures with PIT. Thus the overhead to obtain a single solution is $(1.8N+4.3)$ times the floating point model’s training time, where $N$ is the total number of models trained with PIT, which can be higher than the number of Pareto points. For instance, for GSC, we had to train 4 PIT models to obtain the Pareto front shown in Fig. 6, resulting in a total overhead of 11.5×11.5$\times$11.5 ×. This is approximately 2.7×2.7$\times$2.7 × higher than the total time required by our joint method to obtain the same result.
In the rightmost plot of Figure 5, we report the results on Tiny ImageNet. As observed for the other benchmarks, the different methods lead to similar results for larger DNNs. The benefits of our approach are most visible for smaller models. With a model 12.15% smaller than the w2a8 baseline, we achieve a higher accuracy (62.21% vs. 61.13%). The concatenation of PIT and MixPrec yields slightly better results, with a 10.29% size reduction at iso-accuracy with respect to our solution. With an accuracy of 58.11%, our proposed method is instead able to achieve a 16.96% size reduction with respect to PIT + MixPrec at iso-accuracy. Furthermore, we have an 3.1$\approx$ 3.1$\times$ 3.1 × lower search time overhead (considering the 5 models on the PIT Pareto front).

### v2

@@ Abstract
Deep Learning, Edge Computing, Quantization, Pruning, Neural Architecture Search
{textblock}
[…]
@@ 1 Introduction
Deep neural networks (DNNs) have showcased impressive performance in a wide range of domains; nevertheless, their computational complexity often clashes with the resource limitations of hardware (HW) devices, especially at the edge [1]. Thus, several studies have explored techniques aimed at reducing the computational complexity and memory requirements of DNNs, while preserving their task performance. Pruning[2] and quantization[3] are two of the most popular ones. The former eliminates unimportant computations from a network, reducing the number of parameters and operations, while quantization, in particular with Mixed-Precision Search (MPS), optimizes the data representation of the model’s parameters and activations.
@@
Our code is open-source at: https://github.com/eml-eda/mixprec-pruning. The paper is structured as follows. Sec. 2 covers the required background concepts, and Sec. 3 summarizes the most relevant works on mixed-precision quantization and pruning. In Sec. 4 we illustrate our proposed approach and in Sec. 5 the experimental results. Lastly, Sec. 6 concludes the paper.
@@ 2.1 Quantization and Mixed-Precision Search
Quantization is one of the key optimizations to reduce the complexity of DNNs. In particular, integer quantization is extensively used to replace the floating-point representation of both weights and activations with low bit-width integers. This leads to model compression, and to the usage of integer arithmetic operators, which are faster and more energy-efficient [3]. Moreover, integer quantization also enables the execution of DNNs even on edge devices without a floating point unit. Quantization can be applied on networks already trained in float (i.e., post-training quantization), or its effect can be simulated during training (i.e., quantization-aware training, which is often beneficial for accuracy [3].
@@
Conventional quantization uses fixed-precision, wherein although different parts of a DNN can use different quantization parameters, a uniform bit-width, typically 8 bits, is applied throughout the entire model. Recently, however, mixed-precision strategies have gained attention [7]. They involve using different bit-widths for different sections of a DNN, resulting in additional optimization opportunities in terms of time, memory, and energy consumption, especially when the underlying hardware natively supports operations at sub-byte precision [9, 10]. Nonetheless, determining the optimal allocation of bit-widths to different segments of the network presents a substantial challenge, as it entails exploring a large solution space that grows exponentially with the DNN’s depth. Recently, approaches inspired by Neural Architecture Search (NAS) [16, 17] have been applied to solve this so-called Mixed-Precision Search problem. Such approaches can be classified into two families, namely iterative and one-shot.
@@ 2.2 Pruning
According to the pruning granularity, two main approaches can be identified, i.e., unstructured and structured. Unstructured pruning considers the removal of single weights. This approach leads to the largest theoretical reduction of parameters and operations without accuracy loss, reaching values as high as 90% [22]. However, it leads to sparsely connected networks that cannot be easily accelerated in hardware, especially on general-purpose platforms, without adding a significant overhead [23, 21]. In fact, sparse computations make it complex to achieve high utilization on parallel hardware and damage the regularity and locality of memory access patterns, resulting in worse cache behaviour [23].
@@ 3.1.1 Non-differentiable techniques
Several recent works have been devoted to the exploration of automatic methods to optimally assign different precisions to different parts of DNNs. The first attempts trying to solve the MPS problem considered only task performance as optimization goal. HAWQ-V2 [34] uses a sensitivity-based heuristic approach that considers second-order Hessian information. Lin et al. [35] show an analytical method that optimizes signal-to-quantization-noise-ratio to find the optimal bit-width allocations across the layers of the network. HAQ [18] and ReLeQ [19] propose to use an iterative scheme, where an RL agent drives the bit-width assignment, at the granularity of individual layers, using both task performance and latency or energy measurements from the target hardware as rewards. Such approaches, while effective and flexible, suffer from the problem of long convergence times due to the iterative search space exploration, as discussed in Sec. 2.1.
@@ 3.1.2 Differentiable techniques
One-shot algorithms emerged as a more lightweight solution with respect to the non-differentiable black-box optimization techniques. The earliest methods were based on a supernet scheme, inspired by the Differentiable NAS (DNAS) literature [17]. A multi-path network (the supernet) is constructed, which, in the case of MPS, contains all possible bit-width assignments as alternative paths. The optimization goal then reduces to selecting a single path from this network, and is solved through continuous relaxation using gradient-descent. Wu et al. [27] propose one of the first approaches of this kind, where the optimization is performed considering a cost penalty term associated to the memory along with the standard task-specific loss. Gong et al. [30] propose a similar approach where precisions are explored on MobileNet-V2 architectures using as cost the energy consumption on the BitFusion accelerator.
One of the key limitations of supernet-based MPS is the linear scaling of memory and computational complexity with the size of the search space. That is, adding a new path to the supernet with an operator quantized at a different precision requires duplicating the weights and computations associated with such an operator. For this reason, the search can be performed only on simpler and smaller proxy tasks, severely limiting the applicability of the method. EdMIPS [7] solves this problem by substituting the expensive parallel convolutions of supernet-like approaches with an efficient composite convolution operation. In this approach, a single copy of the float tensors is fake-quantized on-the-fly at the different precisions explored (e.g., 2-, 4- and 8-bit) in each iteration of the gradient-based search.
@@ 3.2.2 Differentiable techniques
DJPQ [32] proposes a joint pruning and quantization scheme where quantization is non-linear and pruning is based on Gaussian gates that control the weight distribution of each layer’s channel. They learn a variational posterior that tries to minimize the mutual information between successive layers’ outputs. Bayesian-bits [31] presents a holistic framework for structured channel pruning and mixed-precision quantization, where pruning is considered as “0-bit” precision. Each precision is associated to a learnable stochastic gate, and the objective function is a balance of task performance and network complexity, measured as bitops. One limitation of this approach is that it is structurally able to explore only precisions which are power-of-two. This can be a drawback for HW platforms such as NE16, which support weights quantized to all possible bit-widths from 2 to 8. Note that this limitation is the mirror image of the one highlighted for FracBits [28]. Moreover, both DJPQ [32] and Bayesian Bits [31] consider a per-layer mixed-precision scheme and rely on cost metrics that poorly correlate with real HW quantities, such as bitops [37].
Chitty-Venkata et al. [33] propose a joint MPS and pruning approach based on gradient-descent. However, quantization is only performed layer-wise and pruning uses a 2:4 block-balanced approach with a fixed sparsity of 50%. This limits the benefits of this approach to hardware platforms that support block-sparse operations. Chitty-Venkata et al. [39] similarly apply layer-wise MPS with a differentiable algorithm, proposing a method to eliminate suboptimal configurations from the search space.
@@ 4.2 Optimization Method
where $W$ is the set of network’s weights, and $\theta:=\{\delta,\gamma\}$ is the set of bit-width selection parameters. The task loss term $\mathcal{L}_{\text{task}}(W,\theta)$ depends on both $W$ and $\theta$. On the other hand, the regularization loss term $\mathcal{R}(\theta)$ is only affected by the bit-width selection parameters, and must be designed to model the desired cost metric, such as size or latency, in a differentiable manner. The relative balance between the two terms is controlled by the scalar strength hyper-parameter $\lambda$. A higher value of $\lambda$ biases the search process towards finding more lightweight networks.
In the search phase, for each weight channel or activation tensor of supported layers (convolutional and linear), an effective tensor is computed, as represented in Fig. 2. More in detail, considering an activation tensor $X^{(n)}$ and its associated bit-width selection parameters $\delta^{(n)}$, the first step consists in obtaining a discrete probability distribution from the latter. Thus, we obtain a new vector $\hat{\delta}^{(n)}$ whose elements are constrained in the $[0,1]$ range and sum up to 1. This can be achieved by applying one among different sampling methods. In this paper we consider three sampling methods, namely a softmax operation (SM), an argmax (AM) operation and a hard Gumbel-Softmax operation (HGSM). We can thus define the sampling operation for the bit-width selection parameters as $h(\delta)$, such that for $i=0,...,|P_{X}|$:
@@ 4.3 Complexity Regularizers
The methodology detailed in Sec. 4.1 and Sec. 4.2 is orthogonal to the cost metric selected as proxy for the DNN’s complexity ($\mathcal{R}$ in Eq. 2). Thus, any measure of DNN cost can be applied to guide the training process, as long as it is expressed by a differentiable function, given that it must be included in a gradient-based optimization loop. Commonly used cost functions are hardware-agnostic proxies for general properties of a DNN, like its size or its number of Multiply-and-Accumulate (MAC) operations [40]. While model size is a good proxy for memory occupation, the number of MACs often correlates poorly with the measured performance (e.g., inference latency or energy consumption) on the real hardware [41]. Therefore, more accurate regularizers can be introduced, which take into account a given platform’s support for each combination of weights and activation precision, and the corresponding efficiency.
@@ 4.3.1 Size Regularizer
where $K_{x}^{(n)}$ and $K_{y}^{(n)}$ are the horizontal and the vertical kernel size. The term $C_{\text{in,eff}}^{(n)}$ represents the effective number of input channels, i.e., those which have not been pruned away. Specifically, for a simple case of sequentially-connected layers, $C_{\text{in,eff}}^{(n)}$ is computed as $C_{\text{out}}^{(n-1)}-\sum_{i=1}^{C_{\text{out}}^{(n-1)}}\hat{\gamma}_{i,p_{0}}$. Using $C_{\text{in,eff}}$ instead of $C_{\text{in}}$ in the cost function models the fact that pruning an output feature map also provides benefits to the subsequent layers since they will process a smaller input.
@@ 4.3.2 Mixed Precision Inference Core Regularizer
To quantify the latency cost of executing a DNN on MPIC, we use the Look-Up Table from [9], which comprises latency measurements collected on fixed layer topologies. In particular, the LUT stores the number of MAC operations per cycle for all combinations of activations and weights bit-widths. For a single convolutional layer $n$, the MPIC regularizer $\mathcal{R}^{(n)}(\hat{\delta}^{(n-1)},\hat{\gamma}^{(n)})$ is defined as:
@@ 4.4 Training Procedure
The procedure to apply the proposed joint mixed-precision quantization and pruning optimizations comprises three distinct phases: warmup, search and fine-tuning. The warmup phase consists in the training of the DNN’s weights only, without the additional bit-width selection parameters. Weights are trained in floating point, and optimized only with respect to the task loss, whereas the regularization loss is not applied. The DNN generated by this preliminary training serves as starting point for the subsequent optimization phases.
@@ 5 Experimental Results
Figure 5: Comparison of the results obtained by our proposed method with other state-of-the-art approaches. The architecture optimized by PIT that is used as input for MixPrec is denoted as PIT Seed.
@@ 5.1 Experimental Setup
We run experiments on three different benchmarks, namely CIFAR-10333http://www.cs.toronto.edu/~kriz/cifar.html, Google Speech Commands (GSC) v2 [11] and Tiny ImageNet [^4] Le, Y., & Yang, X.S. (2015). Tiny ImageNet Visual Recognition Challenge.. For CIFAR-10 we divide the images into training, validation and test sets in the proportions of 66%, 17% and 17%, respectively. The reference architecture is a custom ResNet with 9 convolutional layers, as in [44]. For GSC, as in [11], we consider the standard classification setting comprising only 12 target labels, i.e., 10 keywords from the original dataset and two additional ”Silence” and ”Unknown word” classes. The training, validation and test sets include, respectively, 85%, 10% and 5% of the data samples. The reference architecture is the Depthwise Separable Convolutional Neural Network (DS-CNN) from [44]. For Tiny ImageNet, we divide the original training set into a training and a validation partition with a 90:10 split. Then, we used the official validation partition as test set. The baseline DNN is a ResNet-18. All accuracy results refer to the respective test sets.
@@ 5.2 Sampling Methods Comparison
Fig. 4 reports the results obtained by applying our method on the three benchmarks, using the size regularizer as cost metric. Each point in the curves represents a model trained with a different value of the regularization strength ($\lambda$). Only the DNNs belonging to the Pareto front according to the validation metrics are reported. Each plot shows as baselines the reference architecture trained in floating point (FP) and three fixed-precision versions with the weights quantized at 2, 4, and 8 bits. The activations are quantized at 8 bits since they do not impact the model size. Each color of the curve is associated to a different sampling method.
@@
From these experiments, we conclude that Softmax is the most stable sampling strategy, since it yields the best results on CIFAR-10 and GSC, and has a comparable performance with the other sampling methods on Tiny ImageNet. Thus, we employ Softmax sampling in all following Sections.
@@ 5.3 State-of-the-art Comparison
We present the results on GSC in the central plot of Fig. 5. EdMIPS and MixPrec yield comparable results with each other, with very similar Pareto fronts. Our method achieves an accuracy of 91.60% with a model size of 2.98 kB. At iso-accuracy, we reduce size by 56.17% with respect to the 6.79 kB model found by MixPrec. Our method also leads to comparable results with respect to the concatenation of PIT and MixPrec. This is due to the fact that the reference FP model is highly over-sized, thus the pruning performed by PIT does not remove fundamental channels which cannot be recovered afterwards. Furthermore, quantization at low bit-widths does not wreck accuracy, allowing MixPrec to reduce the precision with negligible performance impact.
However, as shown in Table II, our method significantly reduces the total search time with respect to the concatenation of PIT and MixPrec. Individually, a single PIT epoch requires 1.8×1.8$\times$1.8 × more time than a standard training epoch of the reference DNN, while for both MixPrec and our proposed approach, the overhead is 4.3×4.3$\times$4.3 ×. On the other hand, obtaining a seed for MixPrec when the two methods are applied sequentially requires to first derive the full Pareto curves of architectures with PIT. Thus the overhead to obtain a single solution is $(1.8N+4.3)$ times the floating point model’s training time, where $N$ is the total number of models trained with PIT, which can be higher than the number of Pareto points. For instance, for GSC, we had to train 4 PIT models to obtain the Pareto front shown in Fig. 6, resulting in a total overhead of 11.5×11.5$\times$11.5 ×. This is approximately 2.7×2.7$\times$2.7 × higher than the total time required by our joint method to obtain the same result.
In the rightmost plot of Fig. 5, we report the results on Tiny ImageNet. As observed for the other benchmarks, the different methods lead to similar results for larger DNNs. The benefits of our approach are most visible for smaller models. With a model 12.15% smaller than the w2a8 baseline, we achieve a higher accuracy (62.21% vs. 61.13%). The concatenation of PIT and MixPrec yields slightly better results, with a 10.29% size reduction at iso-accuracy with respect to our solution. With an accuracy of 58.11%, our proposed method is instead able to achieve a 16.96% size reduction with respect to PIT + MixPrec at iso-accuracy. Furthermore, we have a 3.1$\approx$ 3.1$\times$ 3.1 × lower search time overhead (considering the 5 models on the PIT Pareto front).